JEOL JEM-4000FX TEM Archive February 13, 2008 – January 30, 2012

JEOL JEM-4000FX TEM

JEOL JEM-4000FX TEM

January 30, 2012 As you may know, we operate the TEM at 3 different acceleration voltages: 100, 200 & 300kV. For each voltage there is a a set of digital camera preference files that a) provide the correct magnification calibration, and b) the proper gain files. The calibration file is self explanatory. The gain file alters the displayed image to account for artifacts. The artifacts are most often seen at 100kV as white spots in the images, however, the adjustments for low voltages must be changed when the TEM is operated at higher voltage settings.

Until now, I handled the process of copying the proper files to match the operating conditions myself. That has changed. The IT department just wrote small files that automate that process, so you can do it yourself. The files are located on the Desktop and are named:

100kV – Digital Micrograph 200kV – Digital Micrograph 300kV – Digital Micrograph

The files are used as follows:

Select the proper file for the operating conditions you’ve chosen, i.e. for 100kV operation you chose 100kV – Digital Micrograph (DM). Double click the file, after ~5 seconds this message will appear. “Please exit and running instances of Digital Micrograph”. At that point you exit the DM program, note that it takes 10-15 seconds to close the program. Now click the OK button. DM will restart with the new preference files. It is a good idea to use this procedure each time you use the TEM.

I will attach this message to the LCD monitor for your reference.

Let me know if you have questions.

January 16, 2012 TEM is operational and ready to be used.

Note that the TEM course labs begin this week. The TEM will be reserved all day Tuesday and Thursday until the week of April 16th.

January 16, 2012 We have a vacuum problem with the TEM. It is not operational. I’m working on it and I’ll post an update later this afternoon.

December 7, 2011 The EDS system has crashed and is unusable. I’m working on it now.

December 6, 2011 The EDS system is up and running on the TEM. We are able to detect carbon with the new detector window. Also, the energy resolution is improved from 200eV to 140eV.

Many of you already know that crossing a grid bar with the detector in results in a control module lock-up. That is, the system freezes and won’t respond without a power reset. The problem is that, as you cross the grid bar, the count rate goes very high and the system cannot process that many counts. This is a design flaw and cannot be repaired. Because of the increased sensitivity that problem is worse. You must decrease the spot size and/or crank the detector out slightly as you cross grid bars. You can monitor count rate as the dead time increases. You want to be ~50% dead time.

November 29, 2011 There is a problem with the EDS system. I am working to sort it out, but it cannot be used now.

I’ll keep you updated as the repair progresses. Thanks for your patience.

November 29, 2011 I’m having some problems with the EDS calibration. It may require a service call later this week. I’ll keep you posted.

November 27, 2011 I have installed the EDS detector and will make the last adjustments today and tomorrow. The TEM should be ready to use by Tuesday.

November 25, 2011 I am reinstalling the EDS detector over the holidays. I hope to have it back online early this week. The TEM cannot be used until the installation is complete.

October 12, 2011 Delay EDS detector for TEM: As I mentioned before, we are having a thin window installed on the EDS detector. Much to my surprise it will take longer than expected to have that done. At this moment the detector should be returned to us around the 2nd week of November. I’ll send you more information as I get it.

October 4, 2011 MY5250 Introduction to Transmission Electron Microscopy will be offered in the Spring 2012 semester. Successful completion of this course will qualify students for independent use of the microscope. Enrollment will be limited to 12 students.

Please pass this information to students and faculty.

September 28, 2011 TEM x-ray detector: I plan to send the EDS detector on the TEM in for a rebuild next week. It will be in the shop for 2 weeks or less.

Once reconditioned it will have improved energy resolution and a thin window with the capability to detect light elements (carbon or higher Z).

September 14, 2011 You do not need to use liquid nitrogen in the ACD’s if the humidity is less than 50%.

September 7, 2011 The Auto-ramp device brought the filament to saturation in 5 minutes today. You’ll remember that it should take about 15 minutes. Please keep an eye out for this problem. If it keeps happening, I can make an adjustment. Let me know if you see this happen again.

August 22, 2011 The TEM is back online and can be used.

Note that there are vertical lines on the CRT. This is an intermittent problem and will correct itself.

August 19, 2011 The chiller is down in the TEM lab and the microscope is unable to be used. Owen will be back next week to fix it. Please arrange your research accordingly.

July 19, 2011 The TEM can be used, but you will need to use liquid nitrogen in the ACD’s. Only use the righthand ACD. The proper order is to 1) put your sample into the TEM and 2) fill the ACD. Remember to refill the ACD if you have an afternoon appointment. Last user of the day heats the ACD with the ACD2 button.

Let me know if you have trouble.

July 18, 2011 TEM is down – LN2 canceled. The TEM HV went offline this weekend due to the heat/humidity. Outdoor temps will be lower for the next 2 days and, hopefully, I can get it back online. I’lll keep you updated.

I’m canceling the LN2 training this morning.

July 11, 2011 TEM users – liquid nitrogen: The lab temp and humidity level are increasing to the point where we need to begin using liquid nitrogen in the Anti COntamination Devices (ACD). I’d like to meet with you all Monday, July 18 at 10A to review how to do this safely.

Please confirm that you will attend.

July 8, 2011 Data storage policy: The power outage last Friday crashed the Gatan TEM computer. Some data was corrupted when the hard drive was shut down instantly.

This situation can happen at any time, on any of the ACMAL computers, and is the reason why we cannot be responsible for storing your data. It is your responsibility to take data off the system when you complete a session.
July 7, 2011 TEM operational: We have repaired the computer and the TEM can now be used.

Please remember to add the accelerating voltage to you calendar reservations.

July 5, 2011 TEM Gatan PC down: The Gatan (digital camera) PC is inoperable. I have called for help and will update you about the repair progress.

June 2, 2011 Please include the accelerating voltage value you’ll be using when you reserve time on the TEM . If you intend to use 300kV after a 200kV user allow an at least an extra 30 minutes for the HT ramp program to run.

Let me know if this is confusing.

May 24, 2011 I’ve opened Friday afternoon on the TEM calendar for those who attend the Thursday seminars.

May 21, 2011 The TEM seminar is provided for people who are currently using the TEM. I am not making room in the seminar for those who aren’t using the TEM now or who may use it later.

May 20, 2011 This is the schedule. No more space is available.

Thursday 10AM Xiaobao Chee Boyi Zhiyong Hessam Hasti

Thursday 2PM Kristen Taile Qi Aaron Tery Ji Qi

IMPORTANT NOTE. This seminar is to instruct _current_ TEM users how to handle the HV system. If you are not currently using the TEM this seminar is not for you.

May 20, 2011 TEM users – Tuesday seminar is full.

May 20, 2011 The TEM is again operational at 300kV.

The HV tank was damaged when the accelerating voltage was raised to 400kV. The HV tank is not conditioned to operate above 315kV and this damage can happen again if you do not use great care during operation. We were very fortunate to have a spare HV tank that works fine. The potential repair cost could have been $12,000 plus parts. >>>>>Please note that extraordinary repair costs such as this directly lead directly to higher hourly use charges.<<<<<

I will be conducting short seminars in order to ensure that each of you understand the HV system operation. THIS SEMINAR IS MANDATORY! You will not use the TEM until you have attended the seminar. No exceptions. No excuses.

Two seminars are scheduled next week – 10-11AM, Tuesday, May 24th and 2-3PM, Thursday, May 26. I’ll be looking for your replies to this message.

April 20, 2011 TEM users – Urgent please read We are going to have the TEM repaired so that 300kV is again usable. The service visit is tentatively scheduled for the week of May 16. This is a significant repair that may involve a 2nd week of service which is planned for the next week – May 23.

Please let me know right away if this schedule will cause you problems.

April 15, 2011 The TEM login computer intermittently shuts off. I am not sure why but I’m working on it. We will be installing new login system hardware soon and the problem will be moot.

Please be patient.

April 8, 2011 TEM images: Since the TEM was damaged, the beam is “flickering” or unstable – and that effect can be seen in the images under the Focus setting. The effects will show up in the Search setting as out-of-focus images. As I mentioned in an earlier email the solution to this is use a smaller condenser aperture and a smaller objective aperture.

April 4, 2011 I noticed today that the beam is unstable at 200kV – the beam flickers. This is a consequence of the recent damage. You can minimize this by using a small condenser aperture and objective aperture.

We have tentatively setup a service visit to repair the HV tank damage in late April. I expect the TEM will be unavailable at least 1 week.

March 8, 2011 I’ve worked with the service engineer today and we think that the TEM is operational at 200kV. I checked beam stability all day and it appears good.

Do not increase the kV beyond ___200KV___. It will be unnecessary to change the kV under any circumstances. Do not use the HTUP program.

Let me know if you see anything unusual.

March 8, 2011 Serious TEM problem: A component has failed inside the high voltage tank that prevents the TEM from being operated above 100kV. This means that high-resolution imaging is impossible. Unfortunately, this is not a problem I can repair and I expect it to take 3-4 weeks before our service engineer can come here to make the repair.

I’ll keep you informed.

February 7, 2011 You will be using the TEM film camera in your next lab. I’ve attached instructions.

January 31, 2011 I’ve a attached a revised ABBREVIATED operating manual. We will be using the ABBREVIATED manual version this week. Please make yourself acquainted with this document before lab. Transfer any notes to this version. Again, remember, the goal is that you won’t need to read every word of the document as you align and operate the TEM. Consider that you cannot read each instruction and complete the practical exam final in the allotted time. Spend whatever time you require to become very familiar with the instructions, especially location of the knobs and switches. The images I sent you 2 weeks ago will help with this. You may sit at the TEM and locate knobs/switches when no one is using the TEM. Let me know if you have any questions.

January 31, 2011 Some of you have noticed that the vacuum drops when using the ACDs. One of them has a vacuum leak. I will be testing to locate the defective ACD but, don’t use them in the meantime. Once I locate the defective ACD, we can use the other, good one.

I’ll put a note on the TEM about the failed unit.

January 26, 2011 I’ve a attached a revised full operating manual. Please transfer your notes to this version and discard the old one. We will begin using the abbreviated operating instructions next week so transfer any notes to it that you will need. Remember, the goal is that you won’t need to read every word of the document as you align and operate the TEM. Consider that you cannot read each instruction and complete the practical exam final in the allotted time. Spend whatever time you require to become very familiar with the instructions, especially location of the knobs and switches. Let me know if you have any questions.

January 13, 2011 We have opened an additional lab section for the TEM course this year at 2-5 PM, Tuesday afternoons. CRN 14770.

Be aware that we offer the TEM course once each year.

I think enrollment closes tomorrow.

January 5, 2011 EDS systems are working. Sorry for the delay.

December 20, 2010 The EDS system on FESEM, 6400 SEM and TEM will be offline beginning tomorrow afternoon. Offline means they won’t work and cannot be used.

I will bring them back online Monday, Jan 3rd.

November 10, 2010 Some of you, who operate the TEM at 300kV, have experienced the beam problem between 60 and 80kX mag. I haven’t been able to repair the problem but I have found a way around it.

In the LH drawer there is a long row of buttons that turn the various lens on and off. One of them is named “CM”. If you are operating at 300kV turn that lens off (lamp will be dim when it is off). Don’t forget to turn it back on when you are done. Scope users at other kV’s should check to make sure the lens is on (button lit) before operation.

I will continue to search for the permanent solution to the problem.

Thanks for your patience.

November 5, 2010 1. We are having a intermittent problem with the beam alignment at 300kV (not 100 or 200) between 60 kX and 80kX. When you increase the magnification from 60 to 80 the beam will become mis-shaped into a “bowtie” or “butterfly” shape. If this happens to you, call me and I can fix the problem. I’m working with the JEOL people to determine the cause.

2. When you are finishing a session, please check the schedule to see if another user is coming in after you. If so, do not install the ACD heaters. Instead, refill the ACD’s for the next person.

October 14, 2010 I will bake the TEM column this weekend. The scope won’t be available from Friday afternoon until Monday ay ~noon.

September 13, 2010 The Autoramp immediately saturated the filament today when it should have taken about 15-20 minutes. I tested it a second time and it worked fine, but I want you all to turn the Filament knob back to ~5 when you are using the Autoramp. When the Autoramp LED is solid green then slowly bring knob to the stop (~6).

Also, the film camera isn’t working. I’ll look into it tomorrow.

It would be great if some of you started using the scope at 300kV ….

September 7, 2010 We will likely be taking on a special FIB/TEM project late this month or early October. I’ll need 3 continuous days of FIB and 2 days of TEM.

Please let me know ASAP if this will cause you problems.

August 30, 2010 The TEM will now operate at 300kV. We should see better resolution and the 300kV beam will penetrate thicker specimens. In order to maintain that capability we will have to slightly change the operating procedures. I’ve attached a draft document outlining the changes.

Additionally, the STEM (Scanning Transmission Electron Microscope) system works now. I need to make a few calibrations. The STEM system will permit secondary (SE) and transmitted images as well as x-ray maps and x-ray line scans. I will begin to develop an instruction manual for the system.

New_instructions_for_TEM_300kV_operation2

August 13, 2010 The TEM will be down for maintenance the entire week of August 23rd. Plan accordingly.

July 20, 2010 A TEM maintenance visit is scheduled for the week of August 23rd and I will be here but unavailable during that week.

June 14, 2010 We have had yet another hard drive failure in the Gatan computer. We are working to get it back online as soon as possible. Probably won’t be repaired today.

The drives don’t appear to be dependable so I hope you are all saving your data off line.

June 1, 2010 I found the specimen holder leaking air into the column this morning. It had a very big fiber across the o-ring. It is easy to check with the optical microscope. Please check the sample holder o-rings for fibers, etc.

May 26, 2010 I mentioned in an email that we had begun to use the ACD’s this week. I thought you all remembered how to use them from last year. MAybe not – the dewars warmed up during a session and the scope shut down. The notes below are a provided as a refresher. Again, let me know if you have trouble.

These are liquid nitrogen cold traps that scavenge contamination from the specimen chamber. The procedure is briefly described here. Fill silver dewar with LN2 from Rm 637. Column vacuum must be 1 x 10-3 Pa or lower. ACD’s are located on either side and slightly behind the column somewhat above the specimen chamber. Cover the viewing port. Roll the stairs to the front of the TEM. Remove heaters or plug from RH ACD inlet. Insert special funnel and slowly add LN2 from silver dewar until full. Repeat on other side. Note that the ACD is harder to reach in LH side due to location of EDS detector. Wait 15 minutes and refill. Cover inlet ports with white caps. Wait 15 minutes until vacuum improves. Under most conditions the column vacuum will reach mid 10-6 Pa range.

When your session ends, you have to boil off the LN2 using the ACD heaters – ESPECIALLY IF YOU INTEND TO REMOVE THE HOLDER.. If you dump the column vacuum water vapor can condense on the ACD plates in the chamber and make a BIG problem. At the end of your session, remove the white caps and replace with the heaters. Connect the electric cable. In the LH drawer, press both red ACD buttons. This turns off the column ion pump. The ACD’s can be turned off after 1/2 hour by depressing and holding the ACD buttons down until the button lights go off. Note that the column vacuum will still be too poor to use the scope.

We will be happy to show you how to do these steps.
May 25, 2010 We have started to use the ACD’s since the humidity is so high. You will have to allow an extra 1/2 or more to get the column vacuum into the mid 10^-5 range. Please let me know if you are not familiar with using the ACD’s!

May 21, 2010 We will have 1 week of maintenance on the TEM, probably the week of June 28. Plan your work accordingly.

January 21, 2010 We have installed 4pi hardware and Revolution software on the FESEM, 6400 SEM and TEM. Instructions for using the software can be found herehttp://electronmicroscopy.org/ Navigate to Hitachi S-4700 FE-SEM > Microanalysis. Note that the instructions are written for the FESEM and some microscope operations will be different. We are actively working on revised instructions for the 6400 and TEM. They should be published within a week or so. Few of you use standards-based quantitative microanalysis and instead use semi-quantitative analysis. I’ve mentioned in earlier messages that depending on the geometry and composition of your sample, error in the results could approach 20%. I think it would be a good idea to use an off-line microanalysis software package to verify your results. Please download that software, DTSA II, here. http://www.cstl.nist.gov/div837/837.02/epq/dtsa2/index.html This is Java software and like ImageJ is free (actually, you paid for it through federal taxes). Let me know if you have questions.

January 11, 2010 The TEM is back online. The computer was repaired….again.

January 7, 2010 The 6400 SEM is working with new 4 pi software for EDS (same as FESEM and TEM) and a new revision of dPict (image acquisition). If possible, wait until next week to use it when I can show you how. The TEM is still down due to computer problems. Parts are on order.

January 4, 2010 The TEM computer is broken again. Its going to the shop tomorrow.

December 21, 2009 I am working to get the TEM operable at 300kV. It will take some time but I believe I can make the scope available for 200kV use during the time I am working. Here’s how I see this working. I will continue increasing the kV during the break. Email me if you need to use the scope. The TEM won’t be available “in case you need it”. Only firm appointments will be honored. During those appointments, I will temporarily lower the kV back to 200 for use and afterwards immediately continue towards the 300kV goal. That means the scope will normally NOT be at 200kV. Please bear with me, this is the only time we will have to accomplish the higher kV.

December 17, 2009 4000FX TEM The computer is back (again) and seems to be working fine.

December 15, 2009 The TEM is down with computer problems. Please look for an email update tomorrow. We’ll reschedule any missed appointments.

December 4, 2009 I will be installing a new ion pump on the TEM over the break. The TEM will be offline Dec 21 & 22.

September 25, 2009 We have begun to always use the ACD’s (anti-contaminator devices) on the TEM now. These are liquid nitrogen cold traps that scavenge contamination from the specimen chamber. The procedure is briefly described here

Fill silver dewar with LN2 from Rm 637. Column vacuum must be 1 x 10-3 Pa or lower. ACD’s are located on either side and slightly behind the column somewhat above the specimen chamber. Cover the viewing port. Roll the stairs to the front of the TEM. Remove heaters or plug from RH ACD inlet. Insert special funnel and slowly add LN2 from silver dewar until full. Repeat on other side. Note that the ACD is harder to reach in LH side due to location of EDS detector. Wait 15 minutes and refill. Cover inlet ports with white caps. Wait 15 minutes until vacuum improves. Under most conditions the column vacuum will reach mid 10-6 Pa range.

The dewars must be refilled every 3-4 hours. This is not optional. If you let them warm up the vacuum system will shut down along with the electron beam.

At the end of your session, remove the white caps and replace with the heaters. Connect the electric cable. In the LH drawer, press both red ACD buttons AND Bake.

In the morning the ACD’s may have to be turned off by pushing them until the lamps go off. The Bake will always have to be turned off by simply depressing and releasing the button – no need to hold it down. The column ion pump will come on but will likely shut off again in 2-10 minutes. Toggle the column pump controller off-them back on to restart. You may have to repeat this up to 4-5 times to get it to stay on.

We will be happy to show you how to do these steps.

August 31, 2009 I’ll be starting a column bake today. SHould be back online Wed.

August 18, 2009 When the TEM hard drives failed recently we discovered that some data was stored on those drives but no where else. I failed to make clear to the TEM users that I cannot be responsible for data storage. For the record – your data is your responsibility.

August 18, 2009 The TEM camera and EDS are up now.

August 14, 2009 The TEM EDS is working. Gatan camera is not working but may be by Monday. I’ll advise.

August 14, 2009 The TEM computer is back but not yet on-line. I am working on it.

August 12, 2009 2 of the 3 drives in the TEM computer have failed. They will try once more to recover the files but even if they are successful the directory structure will be lost. We will have to locate files in 300GB of information. In even the best situation we will have to reload all the programs.

August 10, 2009 The TEM is computer is still in the shop and the system is inoperable. The hard drive repair is not going well. The repair folks are trying to save the data on the damaged drive. I’ll get a update from them late this afternoon.

August 6, 2009 We’ve been advised to use the anti-contamination devices (ACD’s), which are liquid nitrogen cold fingers that sit above and below the specimen. Any gaseous contamination in the specimen area is trapped on these cold surfaces.

The abbreviated procedure is to fill the 2 dewars with liquid nitrogen ever 4 hours. When the session is complete the user will connect the heaters to the dewars and push the ACD Heat buttons in the LH drawer. The heaters will evaporate the LN2 and bring the fingers to room temp and then shut off in about 1 hour. During the heating cycle the column ion pump will be shut off and the column will be pumped with only the diffusion pump. When the cycle is complete the system will automatically turn on the ion pump.

And this is the problem. The ion pump does not come back on. It has to be manually toggled back on. I am working with JEOL to identify the issue. It could be that we haven’t used the ACDs in so long that there is a tremendous amount of contamination that is released each time we heat. If that is the case the problem will solve itself over time.

In the meantime, we will continue to use the ACD’s. If you need help please let me know and we’ll do it for you.

I will keep you informed.

July 28, 2009 The vacuum recovery after inserting a holder was taking over 6 hours. I baked the column ion pump and lens over the weekend. The column vac is 2X10-5, a considerable improvement. I haven’t evaluated pumping speed after holder insertion.

June 29, 2009 Water vapor is getting into the column when air is accidentally admitted during sample insertion. The water molecules are slow to pump off the inner surfaces. Two things will happen.

I will bake the column. We will quit accidentally admitting air. To accomplish this only Anjana and I will insert the samples.

The scope can be used now.

June 8, 2009 The HV shuts off intermittently but only when the scope is not being used. This is a problem I’ve addressed before and I think can be remedied with a small adjustment.

April 15, 2009 The LH goniometer control “slips” when turning in the CW direction. The problem can be overcome using the 1-2 position control. I will save the repair until we vent the column for something else.

April 2, 2009 The new 4pi hardware was successfully installed today.

February 11, 2009 The TEM is ready to use. I have replaced the broken goniometer with a spare and replaced the condenser apertures. The scope is aligned at 200kV.

Note about the goniometer, there is a limited range that the stage can be moved. If you feel any resistance at all, stop turning. If you think there is something wrong, that is, the stage should more farther, call me to check the problem right away.

February 11, 2009 I have replaced the goniometer with a spare and replaced the condenser lens apertures. The scope is working and can be used.

February 2, 2009 The TEM stage drive is broken and the scope cannot be used. I’ll update you when it’s back online.

January 26, 2009 If you are a TEM users, read on.

I’ve aligned the TEM at 100kV and 200kV. I ask that you always return the TEM to 200kV.

Here’s how to change the kV.

1. Ensure that the filament is unsaturated (red LED on Autoramp).

2. The airlock must be closed.

3. Use the toggle switch above the HT button, labeled 4 on the attached pdf file.

4. Turn up the CRT intensity on the TEM (left hand side screen)

5. Toggle the switch up to raise the kV and down to lower. Watch the kV number on the CRT as you toggle the switch.

JEOL_JEM-4000FX_TEM_change_kV

Please let me know if you have any questions.

January 23, 2009 TEM of nanotubes

I saw this message on the Microscopy Society of America listserver. Maybe some of you would like to join the listserve? Follow this link http://www.microscopy.org/ and look for “Microscopy Listserver” under “Reference and Education Activities” section.

January 22, 2009 Always use the booking calendars to make appointments to use the scopes – even for scopes that don’t get much use like the TEM. I need to see who is using the scope and when.

October 29, 2008 The column vacuum drops badly when the stage position control is moved. The vacuum drop is enough to shut off the HV. YOu can prevent this by simply moving that control BEFORE you open the airlock. I’ll look into a fix. Be on the lookout for things like this to come up since we are using the TEM far more than even a few months ago.

October 12, 2008 The stage pre-pump is not working right and the column vacuum drops when the stage is inserted. I’m investigating.

October 2, 2008 We installed and trained with the in-situ AFM stage and controller. Please use the calendar to reserve time since use will increase.

July 17, 2008 We changed the recirculator water and filters.. Ethylene glycol was used to prevent algae growth. After the shutdown the lens water control valve failed. I removed the valve guts and the system works fine.

We continue to observe that the accelerating voltage will shut off for some reason. I am still troubleshooting that problem but the scope is usable.

The junk is still visible in the digital images. Not film images.

February 13, 2008

4000 TEM, beam is aligned. There is some object on the digital camera screen.